OpenStudy (anonymous):

can you please help me with my home work?? I need some answers fast! 2{8y-[4-7(3y-5)]}

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

2(8y-(4-21y+35)) 2(8y-(39-21y)) 2(8y-39+21y) 2(29y-39) 58y-78

OpenStudy (anonymous):

good answer!!! Wait I have more!!

OpenStudy (anonymous):

1/4(28y-4)-1/3(-12y-9)-13

OpenStudy (anonymous):

what is 4 there? is it 1/4 of the (28y-4)?

OpenStudy (anonymous):

sorry it is a fraction 1/4. the homework I am doing is asking to collect like terms...

OpenStudy (anonymous):

okay, 1/4(4(7y-1))-1/3(3(-4y-3))-13, cancel 4 and 3 (7y-1)-(-4y-3)-13 7y-1+4y+3-13 11y-11

OpenStudy (anonymous):

or we can have 11(y-1)

OpenStudy (anonymous):

thanks dude you are awsome!!

OpenStudy (anonymous):

You are welcome. Don't I deserve some medals?hehehe,,lol

OpenStudy (anonymous):

Thank you for the compliment!

OpenStudy (anonymous):

hey guys need a little help!!! (1/4,-1/2) and (3/4,5) this is asking me to find an equation of the line containing the given pair of points. please help!!

OpenStudy (anonymous):

hey guys need a little help!!! (1/4,-1/2) and (3/4,5) this is asking me to find an equation of the line containing the given pair of points. please help!!

OpenStudy (anonymous):

what are the set builder numbers greater than or less than 44

OpenStudy (anonymous):

use this formula (y-y1)=(y2-y1)/(x2-x1) (x-x1) (1/4,-1/2) and (3/4,5) y-5=(5+1/2)/(3/4-1/4) (x-3/4) y-5=(11/2)/(2/4)(x-3/4) y-5=11(x-3/4) y-5=11x-33/4 y-11x=5-33/4 y-11x=-13/4

OpenStudy (anonymous):

thanks man!! How about -(2)/(13)*(1)/(2)*(-(26)/(15))

OpenStudy (anonymous):

just multiply the numerators (-2, 1 and -26) the answer is 52 multiply the denominator, (13,2 and 15) the answer is 390. so, the fraction is 52/390 or 2/15

OpenStudy (anonymous):

hey man you are greatttt! what about -1/y-y and what about -4/2-2

OpenStudy (anonymous):

-1/y -y, the LCD is y so it will become, (-1-y^2)/y

OpenStudy (anonymous):

-4/2 - 2, as you can observe -4/2 is -2 so this will become 2-2=-4

OpenStudy (anonymous):

thanks man you are a big help!!! I need so much help with my homework til it's not funny!!! Are you usually on at this time?? because I have so much more homework to do...

OpenStudy (anonymous):

do you have skype?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

because i can explain it well on skype

OpenStudy (anonymous):

okay i will help

OpenStudy (anonymous):

thanks a lot I appreciate that!!